OpenStudy (anonymous):

find the mid point of a segment with the given end points A(3,-5)and C(2,9)

jimthompson5910 (jim_thompson5910):

to find the midpoint, you average the x coordinates to get (x1+x2)/2 = (3+2)/2 = 5/2 = 2.5 and you do the same for the y coordinates (y1+y2)/2 = (-5+9)/2 = -4/2 = -2 so the midpoint is (2.5, -2)

OpenStudy (anonymous):

thanks:)

jimthompson5910 (jim_thompson5910):

yw